OpenStudy (anonymous):

A ball is dropped from the top of a building that is 250 feet tall. The height h of the ball in feet after t seconds is modeled by the function h = -16t2 +250. How long will it take for the ball to reach the ground? Round your answer to the nearest hundredth if necessary.

OpenStudy (anonymous):

h(t) = 0 0 = -16t^2 + 250 16t^2 = 250 /16 /16 t^2 = 15.625 square root t and 15.625 T= 3.953 Seconds

OpenStudy (anonymous):

how do i verify my answer?

OpenStudy (anonymous):

Plug your answer (3.953) back into the equation and solve for H(t) to verify that it comes out to 0.

OpenStudy (anonymous):

Let me know if you need it worked out, but it seems like you got the hang of it.

OpenStudy (anonymous):

I need it worked out pleeasse

OpenStudy (anonymous):